Circular No. 8919

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2007 V11–V14, C/2007 W4–W12, C/2007 X1–X4 (SOHO)

Further to IAUC 8918, additional near-sun comets have been found on SOHO website images, as tabulated below; all were Kreutz sungrazers except for C/2007 X1 (no known group affiliation). C/2007 V11 was bright (mag 4) with a short tail in C3 images; in C2 images, it also showed a short, faint tail. C/2007 V12 was diffuse with a hint of faint tail (mag ∼ 7). C/2007 V13 and C/2007 X2 were extremely faint (mag ∼ 8.5) and diffuse. C/2007 V14 was tiny and stellar in appearance (mag ∼ 7.5). C/2007 W4 and C/2007 X1 were small and stellar in appearance (mag ∼ 6.5). C/2007 W5 was tiny, very faint (mag ∼ 8), and slightly diffuse. C/2007 W6 was tiny, very faint (mag ∼ 7.5), and quite diffuse. C/2007 W7 was stellar in appearance (mag 6.5) in C3 images, and it appeared as a bright ‘teardrop’ in C2 images. C/2007 W8 was faint (mag 7.5) and very diffuse. C/2007 W9 was very small and diffuse (mag 8). C/2007 W10, which was also found by M. Kusiak, was very diffuse and elongated (mag 7.5). C/2007 W11 and W12 were small and somewhat diffuse (mag 7). C/2007 X3 was tiny, faint (mag 7.5), and slightly diffuse. C/2007 X4 was stellar in appearance (mag ∼ 5.5) in C3 images, and quite condensed and teardrop-shaped in C2 images.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 V11 Nov. 10.596 14^{h}38.^{m}9 −21^{o}10′ C3/2 HS 2008-B09
C/2007 V12  11.118 14 59.6 −18 41 C2  SY 2008-B09
C/2007 V13  13.435 15 09.3 −19 26 C2  RK 2008-B09
C/2007 V14  14.643 15 14.5 −19 44 C2  RK 2008-B09
C/2007 W4  19.113 15 27.2 −22 21 C3  MA 2008-B10
C/2007 W5  19.310 15 34.0 −20 57 C2  HS 2008-B10
C/2007 W6  22.758 15 49.2 −21 48 C2  RK 2008-B10
C/2007 W7  23.154 15 42.3 −24 10 C3/2 BZ 2008-B10
C/2007 W8  23.692 15 53.5 −22 03 C2  HS 2008-B10
C/2007 W9  23.217 15 54.3 −22 09 C2  TH 2008-B35
C/2007 W10  24.125 15 55.9 −22 11 C2  HS 2008-B35
C/2007 W11  29.354 16 20.2 −23 20 C2  RK 2008-B35
C/2007 W12  30.829 16 27.0 −23 34 C2  HS 2008-B35
C/2007 X1  Dec. 3.064 16 29.0 −21 01 C2  MU 2008-B36
C/2007 X2  4.713 16 44.9 −24 07 C2  HS 2008-B36
C/2007 X3  4.814 16 44.2 −24 07 C2  RK 2008-B36
C/2007 X4  5.388 16 41.8 −26 35 C3/2 HS 2008-B36
2008 February 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green